A company makes electric motors. The probability an electric motor is malfunctioning is 0.01. What is the probability that a sample of 300 electric motors will contain exactly 5 defective motors?


1
Expert's answer
2021-03-09T02:26:05-0500

We use Poisson's formula:

"{P_n}(k) = \frac{{{\lambda ^k}}}{{k!}}{e^{ - \lambda }}"

"\lambda = np = 300 \cdot 0,01 = 3"

Then

"{P_{300}}(5) = \frac{{{3^5}}}{{5!}}{e^{ - 3}} \approx 0.1"

Answer: 0.1